OpenStudy (anonymous):

10c-f=-13=cd solve for c. show work please

OpenStudy (anonymous):

10c - f = -13 = cd set up 2 equations 10c - f = -13 and -13 = cd solve one at a time: 10c - f = -13 10c = -13 +f c = (-13 + f) / 10 -13 = cd -13/d = c

OpenStudy (anonymous):

c = (-13 + f) / 10 and c = -13/d so it can also be said that (-13 + f) / 10 = -13/d